The greenhouse floor has a diagonal equal to 28 feet. The sides are square. What is the length of one side?
Answer by ptaylor(2198)

The greenhouse floor has a diagonal equal to 28 feet. The sides are square. What is the length of one side?
We'll use the Pythagorean Theorem:
28^2=s^2+s^2
28^2=2s^2 take sqrt of both sides:
28=s(sqrt2)

s=28/sqrt2=19.799 feet
ck
(19.799)^2+(19.799)^2=28^2
392.005+392.005=784
784.010~~784

Hope this helps---ptaylor